OpenStudy (anonymous):

Consider this equation: 4x – 5 + 6x = 6 Which is an equivalent equation after combining like terms?

OpenStudy (anonymous):

Consider this equation: 4x – 5 + 6x = 6 Which is an equivalent equation after combining like terms? A. -1 + 6x = 6 B. 10x - 5 = 6 C. 4x + 1 = 6 D. 4x - 5 = 0

pooja195 (pooja195):

4+6=?

OpenStudy (anonymous):

10

pooja195 (pooja195):

so 10x which is yr answer

OpenStudy (anonymous):

For which equation is x = 9 a solution? A. 3x – 7 = –34 B. x – 9 = 1 C. 2003-05-02-00-00_files/i0020000.jpg D. 4x + 5 = 41

pooja195 (pooja195):

Welcome to openstudy! One question per post please ! :)